OpenStudy (anonymous):

Help please Determine the direction that this parabola opens y=2x^2-4x+5 Up or down?

OpenStudy (vishweshshrimali5):

Hi there! It's an up parabola

OpenStudy (anonymous):

How do you know though?

OpenStudy (vishweshshrimali5):

I can write the equation as: \[y = 2(x^2 - 2x + 1) + 3 = 2 (x-1)^2 + 3\] Now, look closely on the RHS: \((x-1)^2\) will always be positive As the value of x increases, y also increases. That's why, you have an up parabola

OpenStudy (anonymous):

Ah okay. Thank-you very much

OpenStudy (vishweshshrimali5):

Your welcome! Cheers!

OpenStudy (anonymous):

So would y=-2x^2-8x-15 be down?

OpenStudy (vishweshshrimali5):

You got it right!

OpenStudy (anonymous):

Thank-you for helping me understand

OpenStudy (vishweshshrimali5):

(y) Your welcome :)